1. What is the new [H2SO4], if 350.0mL of 6.00M H2SO4 is diluted with 570.0mL of water?


2.What is the new MnSO4 concentration, if 25.0mL of 1.45M solution is diluted to 1.00L?
1
Expert's answer
2020-06-15T14:37:46-0400

Question 1

When a substance is diluted the number of moles of the solute remains the same..Therefore, the following expression is used

"C"1"V"2 "=C"2"V"2

Thus,

"C2=\dfrac{C1V1}{V2}"


"=\dfrac{6.mol.L^{-1}*350mL*10^{-3}.L.mL^{-1}}{570 mL*10^{-3}.L.mL^{-1}}"


"C2" "=3.68M"


Question 2

We use the following expression

"C1V1=C2V2"

Thus

,"C2=\dfrac{C1V1}{V2}"


"=\dfrac{1.4M*0.025L}{1.0.L}"


"C2=0.035M"